Two particles 1 and 2 moves with constant velocities V1and V2. At the initial moment their radius vectors are equal to r1 and r2. how must these four vectors be interrelated for the particles to collide.

Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear biswas,

Answer to your question is (r1-r2) / |r1-r2| = (v1-v2) / |v1-v2|

Here relative velocity concept will be applicable...

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

Sagar Singh

B.Tech IIT Delhi

Shivam
6 years ago
Here,velocity of 2 with respect to 1=v2vector -v1vector/|v2-v1|=r1-r2/|r1-r2|...................................................